OpenStudy (anonymous1):

I have a question about a derivative of trigonometric functions. Details in the reply.

OpenStudy (anonymous1):

I must show that \[\frac{d(\ln \sin x)}{d(\ln x)} = x\tan x\] I've tried it as follows: \[d(\ln \sin x) = x\tan x\ d(\ln x)\] But \[ d(\ln x) = \frac{1}{x}dx \] So \[d(\ln \sin x) = x\tan x\ \frac{1}{x}dx\] \[d(\ln \sin x) = \tan x\ dx\] which is certainly not true. What am I doing wrong?

OpenStudy (anonymous):

U substitution

OpenStudy (anonymous):

they should cancel out after u derive the u subtitution, and u will have an easy derivative

OpenStudy (anonymous):

for the ln(sinx) u will have cosx/sinx as the answer...need more help?

OpenStudy (anonymous):

which is = to 1/tanx and lnx= 1/x, so the equation is equal to xtanx

OpenStudy (anonymous1):

What is wrong with the steps I used above?

OpenStudy (anonymous):

d/dx ( ln(sinx)) = 1/tanx

OpenStudy (anonymous1):

I still don't see it correctly.

OpenStudy (anonymous1):

It also appears to me that x tan x isn't correct. Maybe it is just an error of the book.

OpenStudy (jamesj):

Corrected: d(ln sin x)/d(ln x) = d(ln sin x)/dx . dx/d(ln x) = d(ln sin x)/dx . [ d(ln x)/dx ]^-1 = cot x / (1/x) = x / tan x , etc. etc. I'm confident your proposed answer of x tan x isn't correct.